OpenStudy (anonymous):

The slope of the line passing through P(1,4) and Q(x,1) is -3/5. What is the value of x? Select one: a. 6 b. 7 c. 9 d. 10

jigglypuff314 (jigglypuff314):

\[slope = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \rightarrow \frac{ 4-1 }{ 1-x }=\frac{ -3 }{ 5 }\] so solve for x by either cross multiplying or by first multiplying both sides by 5 then by (1-x) so 5(4-1) = -3(1-x) 5(3) = -3 + 3x 15 = -3 + 3x then add 3 on both sides +3 +3 18 = 0 + 3x then divide both sides by 3 so /3 /3 x = 18/3 which you can reduce down

OpenStudy (anonymous):

6

jigglypuff314 (jigglypuff314):

yes :)